OpenStudy (anonymous):

Solve for x : 8 x + 10 - 5 x = 15.

OpenStudy (anonymous):

i would start by adding like terms

OpenStudy (anonymous):

so add 8x -5x + 10 =15

OpenStudy (anonymous):

3x + 10 =15

OpenStudy (anonymous):

can you do it from here?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

3/5

OpenStudy (anonymous):

okay, so you have 3x +10 =15 you subtract 10 from both sides you get 3x = 5 then divide 3 from both sides and you get...

OpenStudy (anonymous):

1 2/3?

OpenStudy (anonymous):

x = \[\frac{ 5 }{ 3 }\]

OpenStudy (anonymous):

ohhh thanks i get it

OpenStudy (anonymous):